# Stealing Training Graphs from Graph Neural Networks

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have shown promising results in modeling graphs in various tasks. The training of GNNs, especially on specialized tasks such as bioinformatics, demands extensive expert annotations, which are expensive and usually contain sensitive information of data providers. The trained GNN models are often shared for deployment in the real world. As neural networks can memorize the training samples, the model parameters of GNNs have a high risk of leaking private training data. Our theoretical analysis shows the strong connections between trained GNN parameters and the training graphs used, confirming the training graph leakage issue. However, explorations into training data leakage from trained GNNs are rather limited. Therefore, we investigate a novel problem of stealing graphs from trained GNNs. To obtain high-quality graphs that resemble the target training set, a graph diffusion model with diffusion noise optimization is deployed as a graph generator. Furthermore, we propose a selection method that effectively leverages GNN model parameters to identify training graphs from samples generated by the graph diffusion model. Extensive experiments on real-world datasets demonstrate the effectiveness of the proposed framework in stealing training graphs from the trained GNN.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph-structured data pervade numerous real-world applications such as social networks (Hamilton et al., 2017), finance systems (Wang et al., 2019), and molecular graphs (Wang et al., 2022). Graph Neural Networks (GNNs) have shown promising results in modeling graphs by adopting a message passing scheme (Kipf and Welling, 2017; Xu et al., 2018; Wang et al., 2024), which updates a node’s representation by aggregating information from its neighbors. The learned representation can preserve both node attributes and local graph structural information, facilitating various tasks, such as node classification  and graph classifications (Errica et al., 2019).
@@ 2. Related Work
Graph Neural Networks. Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Ying et al., 2018; Liu et al., 2022; Javaloy et al., 2023; Lee et al., 2024; Xu et al., 2024; Lin et al., 2024b) have shown great power in modeling graph-structured data, which have been deployed to various applications such as social network analysis (Fan et al., 2019; Dai and Wang, 2021a; Zhang et al., 2024b; Lin et al., 2024c), drug discovery (Xu et al., 2023; Dai and Wang, 2021b, 2022b; Dai et al., 2023b; Zhang et al., 2024a) and energy network analysis (Dai and Chen, 2022). The success of GNNs lies in the message-passing mechanism, which iteratively aggregates a node’s neighborhood information to refine the node’s representations. For example, GCN (Kipf and Welling, 2017) combines a node’s neighborhood information by averaging their representations with the target center nodes. To improve the expressivity of GNNs, GIN (Xu et al., 2018) further incorporates a hidden layer in combining the neighbors’ information. Inspired by the success of transformers in modeling image and text (Devlin et al., 2018; Dosovitskiy et al., 2020; Lin et al., 2024a), graph transformer is also proposed (Müller et al., 2023; Rampášek et al., 2022; Zhang et al., 2020; Ying et al., 2021). Generally, graph transformer has a global attention mechanism for graph embedding learning. It shows promising results, especially in molecule property prediction. Despite the great achievements, GNNs could be vulnerable to privacy attacks, which largely constrain their adoption in safety-critical domains such as bioinformatics and financial analysis.
@@
Privacy Attacks on GNNs. The training of GNNs requires a large amount of data. In critical domains such as bioinformatics (Li et al., 2021) and healthcare (Li et al., 2020), sensitive data of users will be collected to train a powerful GNN to facilitate the services. However, recent studies show that privacy attacks can extract various private information from GNN models (Dai et al., 2022, 2024; Olatunji et al., 2021; Duddu et al., 2020; Dai et al., 2023a; Dai and Wang, 2022a), threatening the privacy of users. For example, membership inference attacks (Olatunji et al., 2021; Duddu et al., 2020; He et al., 2021b; Wu et al., 2024) can identify whether a target sample is in the training set. This is achieved by learning a binary classifier on patterns such as posteriors that training and test samples exhibit different distributions. Model extract attacks (Wu et al., 2022) aim to steal the target model by building a model that behaves similarly to the target model. Model inversion attacks, also known as reconstruction attacks, try to infer the private information of the test/training data. For example,  Duddu et al. (Duddu et al., 2020) propose to infer sensitive attributes of users from their node embeddings. The reconstruction of the adjacency matrix from node embeddings is studied in (Zhang et al., 2022). GraphMI (Zhang et al., 2021b) further considers a white-box setting for adjacency matrix reconstruction, where the target GNN model parameters and node features are available for attackers. Our GraphSteal lies in the model inversion attack. However, GraphSteal is inherently different from the aforementioned methods because (i) we focus on a new problem of stealing training graphs from the trained GNN without even partial information of training set; (ii) we propose a novel framework GraphSteal which designs a diffusion noise optimization algorithm and a model parameter-guided graph selection mechanism to reconstruct high-quality training graphs of the target GNN model.

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have shown promising results in modeling graphs in various tasks. The training of GNNs, especially on specialized tasks such as bioinformatics, demands extensive expert annotations, which are expensive and usually contain sensitive information of data providers. The trained GNN models are often shared for deployment in the real world. As neural networks can memorize the training samples, the model parameters of GNNs have a high risk of leaking private training data. Our theoretical analysis shows the strong connections between trained GNN parameters and the training graphs used, confirming the training graph leakage issue. However, explorations into training data leakage from trained GNNs are rather limited. Therefore, we investigate a novel problem of stealing graphs from trained GNNs. To obtain high-quality graphs that resemble the target training set, a graph diffusion model with diffusion noise optimization is deployed as a graph generator. Furthermore, we propose a selection method that effectively leverages GNN model parameters to identify training graphs from samples generated by the graph diffusion model. Extensive experiments on real-world datasets demonstrate the effectiveness of the proposed framework in stealing training graphs from the trained GNN. The code is publicly available at https://github.com/ventr1c/GraphSteal.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709289isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph-structured data pervade numerous real-world applications such as social networks (Hamilton et al., 2017), finance systems (Wang et al., 2019), and molecular graphs (Wang et al., 2022). Graph Neural Networks (GNNs) have shown promising results in modeling graphs by adopting a message passing scheme (Kipf and Welling, 2017; Xu et al., 2018; Wang et al., 2024; Liang et al., 2024b), which updates a node’s representation by aggregating information from its neighbors. The learned representation can preserve both node attributes and local graph structural information, facilitating various tasks, such as node classification  and graph classifications (Errica et al., 2019).
@@ 2. Related Work
Graph Neural Networks. Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Ying et al., 2018; Liu et al., 2022; Javaloy et al., 2023; Lee et al., 2024; Xu et al., 2024; Lin et al., 2024b) have shown great power in modeling graph-structured data, which have been deployed to various applications such as social network analysis (Fan et al., 2019; Dai and Wang, 2021a; Zhang et al., 2024b; Lin et al., 2024c; Liang et al., 2024d), drug discovery (Xu et al., 2023; Dai and Wang, 2021b, 2022b; Dai et al., 2023b; Zhang et al., 2024a) and energy network analysis (Dai and Chen, 2022). The success of GNNs lies in the message-passing mechanism, which iteratively aggregates a node’s neighborhood information to refine the node’s representations. For example, GCN (Kipf and Welling, 2017) combines a node’s neighborhood information by averaging their representations with the target center nodes. To improve the expressivity of GNNs, GIN (Xu et al., 2018) further incorporates a hidden layer in combining the neighbors’ information. Inspired by the success of transformers in modeling image and text (Devlin et al., 2018; Dosovitskiy et al., 2020; Lin et al., 2024a; Fang et al., 2023, 2025), graph transformer is also proposed (Müller et al., 2023; Rampášek et al., 2022; Zhang et al., 2020; Ying et al., 2021; Fang et al., 2024; Liang et al., 2024c). Generally, graph transformer has a global attention mechanism for graph embedding learning. It shows promising results, especially in molecule property prediction. Despite the great achievements, GNNs could be vulnerable to privacy attacks, which largely constrain their adoption in safety-critical domains such as bioinformatics and financial analysis.
@@
Privacy Attacks on GNNs. The training of GNNs requires a large amount of data. In critical domains such as bioinformatics (Li et al., 2021) and healthcare (Li et al., 2020), sensitive data of users will be collected to train a powerful GNN to facilitate the services. However, recent studies show that privacy attacks can extract various private information from GNN models (Dai et al., 2022, 2024; Olatunji et al., 2021; Duddu et al., 2020; Dai et al., 2023a; Dai and Wang, 2022a; Hou et al., 2024; Liang et al., 2024a), threatening the privacy of users. For example, membership inference attacks (Olatunji et al., 2021; Duddu et al., 2020; He et al., 2021b; Wu et al., 2024) can identify whether a target sample is in the training set. This is achieved by learning a binary classifier on patterns such as posteriors that training and test samples exhibit different distributions. Model extract attacks (Wu et al., 2022) aim to steal the target model by building a model that behaves similarly to the target model. Model inversion attacks, also known as reconstruction attacks, try to infer the private information of the test/training data. For example,  Duddu et al. (Duddu et al., 2020) propose to infer sensitive attributes of users from their node embeddings. The reconstruction of the adjacency matrix from node embeddings is studied in (Zhang et al., 2022). GraphMI (Zhang et al., 2021b) further considers a white-box setting for adjacency matrix reconstruction, where the target GNN model parameters and node features are available for attackers. Our GraphSteal lies in the model inversion attack. However, GraphSteal is inherently different from the aforementioned methods because (i) we focus on a new problem of stealing training graphs from the trained GNN without even partial information of training set; (ii) we propose a novel framework GraphSteal which designs a diffusion noise optimization algorithm and a model parameter-guided graph selection mechanism to reconstruct high-quality training graphs of the target GNN model.
@@ 6. Conclusion and FUTURE WORK
Acknowledgements.
This material is based upon work supported by, or in part by, the Army Research Office (ARO) under grant number W911NF-21-1-0198, the Department of Homeland Security (DHS) CINA under grant number E205949D, and the Cisco Faculty Research Award. The findings in this paper do not necessarily reflect the view of the funding agencies.